# organization=Instituto Federal de Educação Ciência e Tecnologia do Ceará (IFCE), Campus Tianguá,addressline=Av Tabelião Luiz Nogueira, s/n - Santo Antônio, city=Tianguá, postcode=62324-075, state=Ceará, country=Brazil

### v3

@@ 2 The holographic principle as a prelude to the emergence of spacetime
In the next section, we shall propose that the evolution of the universe’s holographic complexity by its second law can shed some light on the issue of Hubble tension if one considers the perspective of spacetime as an emergent entity according to the lines introduced in [20, 21].
5 Holographic complexity and the Hubble tension: a possible quantum gravitational standpoint for cosmological observations.
Considering the fundamental aspect of the theory introduced in [20, 21], that the degrees of freedom to the emergence of spacetime consist of quantum correlations only, the concept of quantum complexity must be helpful.
In such a situation, quantum states and unitary transformations would not be thought of as "things" but as patterns of connectivity and transformation within such a structure. Moreover, the growth of complexity would reflect structural change, rather than mere dynamical evolution. In such a context, one would have a way to think about change in a scenario where time is absent.
Taking into account the issue of the emergence of spacetime, one can consider that quantum entanglement carried by the structure of correlations can be used as the key resource for the quantum computational process driving such a phenomenon [62, 63].
In this situation, quantum complexity can be given in terms of the number of quantum correlations effectively used to accomplish spacetime emergence. To obtain such a number, one must fix at first the total number of degrees of freedom in the theory by considering both its UV and IR regimes [71].
Concerning the IR regime, one must put the system into a region with a finite size. In the present scenario, such a region must be large enough to contain the necessary holographic information for the emergence of spacetime.
Particularly, in the $AdS/CFT$ framework, causal diamonds defined on the $AdS$ boundary carry such a necessary amount of information [72, 73, 74], being considered as the natural unit of holography in a region of spacetime [75].
Based on such a fact, we shall take the universe’s causal diamond as the region of finite size used to regularize the IR regime of the present theory for the emergence of spacetime.
In this case, we have that the radius of the universe causal diamond is given by
@@
On the other hand, to regularize the UV regime, one must introduce a cutoff length scale in the theory. In the scenario introduced in [20, 21], it can be provided by the area parameter $\Delta$ defined by the Eq. (2). In this case, one can take $\epsilon=\Delta^{1/2}$ as the necessary UV cutoff.
Bringing all these ingredients together, the holographic complexity necessary to accomplish the emergence of cosmic spacetime can be written as
@@
Considering the relationship between $C_{h}$ and the number of quantum correlations effectively used to the emergence of cosmological spacetime, we shall take:
@@
In the expression above, the exponent $\alpha$ will provide a measure of how hard the task of picking quantum correlations up to accomplish the emergence of spacetime (as $\alpha$ increases, harder becomes the task). As we shall see, the exponent $\alpha$ will also provide the saturation limit of $C_{h}$.
Taking into account the spin network description of the theory introduced in [21], we shall consider that the number $N$ in the Eq. (10) can be written in terms of the amount of spin network external legs puncturing the universe causal diamond horizon [21].
In this case, we shall consider that the spin network’s legs patch quanta of area on the causal diamond horizon in a fractal way, according to the following relation:
$N\sim\Big{(}\frac{A_{\Diamond}}{\Delta}\Big{)}^{\varepsilon},$ (11)
where $A_{\Diamond}=4\pi R_{\Diamond}^{2}$ is the horizon’s area. In the Eq. above, $\varepsilon>0$ can be related to the fractal dimension attributed to the causal diamond horizon by considering its microscopic structure description [76, 77] (See Fig. (1)).
Figure 1: Artistic representation of spin network legs crossing the universe’s causal diamond horizon. Each spin network leg patches a quantum of area which is distributed on the horizon surface in a fractal way, according to Eq. (11). The number of spin network legs crossing the horizon $N$ can also be related to the amount of quantum complexity necessary for the emergence of cosmic spacetime. (Figure generated by CANVA®  with the use of AI.)
By tracing the basic assumptions of our construction, let us now present our results. In this way, from Eqs. (8) and (11), we can find that
@@
which together with Eq. (2) gives
@@
Finally, by using Eqs. (9) and (16), we obtain
$H_{0}^{2}\sim C_{h}^{\gamma}\;.$ (17)
@@
$\gamma=\frac{3+3\varepsilon-4\alpha\varepsilon}{3\alpha\varepsilon}.$ (18)
To proceed, it is interesting to analyze the consistency between our results and the usual form of the AdS/CFT correspondence. Moreover, in the present analysis, we shall consider that there exists an upper bound to the exponent $\alpha$ in Eq (9) to obtain a saturation limit for $C_{h}$, according to its second law.
In this sense, we note from Eqs. (10) and (15), that
@@
To have consistency with the AdS/CFT correspondence, $L$ must depend on a positive power of $N$ in the sense of obtaining a low curvature regime when $N$ is large [71]. In this case, we must have
$02\alpha\varepsilon-3>$ 0 . (20)
The only possible solution to $\gamma$ consistent with the existence of an upper bound to the exponent $\alpha$ is
@@
In fact, such a result together with Eq. (20) provides:
$\varepsilon$ $>$ $1\;;$
$\frac{3}{2\varepsilon}<$ $\alpha$ $\leq\frac{3(\varepsilon+1)}{4\varepsilon},$ (22)
where the upperbound to $\alpha$ is given by
$\alpha=\frac{3(\varepsilon+1)}{4\varepsilon}.$ (23)
@@
Considering the second law of holographic complexity, $H_{0}$ must increase during the evolution of the cosmos till it saturates at a maximal value. It must shed some light on the issue of Hubble tension by tracing a way $H_{0}$ can have different values throughout the universe’s evolution.
@@
In the present paper, we have used the quantum gravity scenario introduced in [20, 21] to approach the important issue of Hubble tension. In this case, it was possible to trace an interesting relationship between $H_{0}$ and the holographic complexity associated with the emergence of cosmological spacetime.
@@
In this case, from the results found out in the present paper, $H_{0}$ must increase with the universe’s holographic complexity during the evolution of the cosmos till it saturates. At this point, such a parameter must reach a constant behavior, recovering the description given by the standard cosmological model. This is in line with several results that have pointed out the possibility that $H_{0}$ must vary throughout the cosmological evolution [17, 18, 19, 78, 79].
In front of such results, the difference between the values obtained to $H_{0}$ when one considers early and late-time measurements must be only a testimony that our universe has evolved from a relatively simple state at its early times to a very complex place containing structures like galaxies, stars, quasars, and black holes. In this sense, it is interesting that black holes have been pointed as the structures that carry the maximal amount of quantum complexity in nature [80].
Such a quantum gravitational reasoning about the Hubble tension can bring us a change of perspective on cosmological observations by considering a quantum gravitational standpoint. In this case, it is interesting to note that, in the context of the theory introduced in [20, 21], there is no role for observers (quantum reference frames) in a quantum gravitational context. This is because quantum reference frames are diluted among the fundamental degrees of freedom of the theory since there is no fundamental distinction between them and quantum correlations in this scenario.
Thus an important change of perspective on the description of the macrocosmos becomes necessary in front of the results found in the present work. One can catch sight of this by noting that $C_{h}^{-\gamma}\sim H_{0}^{-1}$ gives a distance scale in cosmology, corresponding to the observable universe’s radius. Consequently, it is possible to understand the different $H_{0}$ values as a manifestation of the universe’s complexity dependence on distance. In this case, the universe’s complexity must decrease when we consider large distance scales, which can be understood on the basis of an important detail of the theory introduced in the present paper: the spin network description of the microscopic structure of spacetime.
In this way, according to Eqs. (9) and (11), the amount of quantum complexity necessary to build spacetime is given in terms of the number of area quanta patched on the universe’s causal diamond horizon by the spin network legs. In this scenario, the number of quanta in a quantized spacetime must depend on what we observe. Particularly, on the scale of distance we observe, where the number of area quanta must decrease as we increase the distance scale due to coarse-graning effects [81, 82].
@@
In this way, different observers in the universe must correspond only to different measures of the holographic complexity associated with the emergence of spacetime by considering different regions of the cosmos. It can open up the range of investigation in quantum cosmology by revealing a new perspective of the world where quantum complexity assumes a central role.
Future explorations in Astrophysics can be done in this direction by considering, e.g., gravitational waves [83], as future detectors become able to measure the quantum complexity amount carried by this kind of signal.

### v4

@@ 2 The holographic principle as a prelude to the emergence of spacetime
In the next sections, we shall propose that some light can be shed on the issue of the Hubble tension if one makes use of the perspective on the emergence of spacetime from quantum correlations introduced in [22, 23], by considering the second law of holographic complexity as giving the way quantum correlations organize to accomplish such a task.
5 A relationship between $H_{0}$ and holographic complexity in a quantum gravity scenario.
By considering the fundamental aspect of the theory introduced in [22, 23], that the degrees of freedom to the emergence of spacetime consist of quantum correlations only, the concept of quantum complexity must be helpful.
In such a situation, quantum states and unitary transformations would not be thought of as ”things” but as patterns of connectivity and transformation within the structure of quantum correlations, where the growth of complexity would reflect structural change, rather than mere dynamical evolution.
Taking into account the issue of the emergence of spacetime, one can consider that quantum entanglement carried by the structure of correlations can be used as the key resource for the quantum computational process driving such a phenomenon [65, 66]. In this case, quantum complexity can be given in terms of the number of quantum correlations effectively used to drive spacetime emergence, and the way they organize to accomplish such a task.
To obtain such a number of quantum correlations, one must at first fix both the ultraviolet (UV) and infrared (IR) limits of the theory [75]. Concerning the IR limit, one must put the system into a region with a finite size, which must be large enough to accommodate the necessary holographic information for the emergence of spacetime.
Particularly, in the $AdS/CFT$ framework, causal diamonds defined on the $AdS$ boundary carry such a necessary amount of information [76, 77, 78], being considered as the natural unit of holography in a region of spacetime [79]. Based on such a fact, we shall take the universe’s causal diamond as the region of finite size used to regularize the IR regime of the present theory for the emergence of spacetime (see Fig. (1)).
The radius of the universe’s causal diamond is given by
@@
On the other hand, we observe that in order to preserve the universe’s causal diamond as the unit of holography used for the emergence of the AdS spacetime, we need such a region, defined on the AdS boundary, to be free from extra-dimensional effects. In this sense, we must bound the size of the universe’s causal diamond as
$R_{\Diamond}<L,$ (9)
since, for $R_{\Diamond}\geq L$, global extra-dimensional effects become important, and we would probe the full extent of the AdS space, leading to distortions in the boundary-to-bulk correspondence. In this way, the condition (9) above defines the regime in which the holographic dictionary applies, imposing an upper bound on the holographic information available to the emergence of spacetime.
Figure 1: Schematic illustration of the universe’s causal diamond: the overlap between the past and future light cones of an observer’s worldline. It shows the region of spacetime that can both influence and be influenced by an observer.
Now, to regularize the UV regime, one must introduce a cutoff length scale in the theory. In the scenario proposed in [22, 23], the cutoff length scale can be provided by the area parameter $\Delta$ defined by Eq. (3), with $\epsilon=\Delta^{1/2}$ as the necessary UV cutoff.
The introduction of such a UV cutoff is possible if one takes into account the usual LQG description of a quantum surface, where each spin network leg patches a quantum of area on the surface. In the present case, such a quantum surface corresponds to the universe’s causal diamond horizon.
In this way, let us consider a superposition of spin networks describing the universe’s causal diamond, where each spin network state in the superposition is characterized by a specific value of the quantum of area $\Delta$, which defines the measurement scale used to probe the emergent geometry, according to Eq. (3).
In this scenario, if one takes $\bar{a}$ as the average quantum of area defined on such a superposition, the number of spin network legs puncturing the causal diamond horizon can be written as
$N\sim\frac{A_{\Diamond}}{\bar{a}},$ (10)
where $A_{\Diamond}$ is the area of the horizon of the universe’s causal diamond.
Now, let us take $\Delta=(1+\beta)\bar{a}$, with $\beta$ measuring the deviation of the quantum of area $\Delta$, characterizing some spin network state, from $\bar{a}$. In other words, $\beta$ can be seen as a quantum fluctuation, or a quantum noise, in the measurement of the area eigenvalues.
In this case, we can write
$N\sim\frac{(1+\beta)A_{\Diamond}}{\Delta}.$ (11)
The expression above links the microscopic picture from which spacetime must emerge to the current observational universe by connecting the IR and the UV cutoffs. We must note that a similar result can be obtained by considering the existence of residual higher-dimensional effects in the definition of the universe’s causal diamond horizon. It is in agreement with the spirit of the BHHC, where higher-dimensional effects in the bulk are translated as quantum corrections in the four-dimensional gravitational theory induced on the AdS boundary.
By considering the expression (11), for a matter of algebraic convenience, we shall rewrite it as
$N\sim\Big(\frac{A_{\Diamond}}{\Delta}\Big)^{\varepsilon},$ (12)
where the exponent $\varepsilon$ is given by
$\varepsilon=1+\frac{ln[(1+\beta)]}{ln[N/(1+\beta)]}.$ (13)
In the present context, the exponent $\varepsilon$ can be interpreted as an effective (fractal) informational dimension of the surface microstate space, which delimits the amount of quantum information that can be used for the emergence of spacetime, despite the existence of quantum noise. We note that $\varepsilon$ approaches 1 for $\beta\rightarrow 0$, corresponding to the minimal quantum noise in the measurement of the emergent area eigenvalues. Such a limit will be discussed later.
By choosing the IR and UV cutoffs above, a superposition of spin networks describing the universe’s causal diamond will be considered as the unit of holography in the present context, accommodating all the holographic information necessary for the emergence of spacetime, despite the existence of quantum noise.
In such a scenario, the holographic complexity necessary to accomplish the emergence of cosmic spacetime can be written as
@@
Now, by considering the relationship between $C_{h}$ and the number of quantum correlations effectively used in the emergence of cosmological spacetime, we shall take:
@@
In the expression above, the exponent $\alpha$ will quantify how efficiently quantum correlations organize to accomplish the emergence of classical spacetime. As $\alpha$ increases, quantum correlations organize more efficiently, which enables the emergence of more complex geometries. On the other hand, as we shall see, there will be an upper bound to such an efficiency, which provides a saturation limit to the complexity associated with the emergent spacetime.
By tracing the basic assumptions of our construction, let us now present our results. From Eqs. (8) and (12), we can find that
@@
which by considering Eq. (3) gives
@@
Finally, by using Eqs. (14), (20), and (15), we obtain
$H_{0}^{2}=DC_{h}^{\gamma}=DN^{\gamma\alpha}=De^{\alpha\log{N^{\gamma}}}\;,$ (21)
where we have englobed all the constants in the calculations above inside the factor $D$.
@@
$\gamma=\frac{3+3\varepsilon-4\alpha\varepsilon}{3\alpha\varepsilon},$ (22)
which will correspond to a cosmological scaling to quantum complexity.
Before proceeding with the analysis of the result above for the exponent $\gamma$, we note first that from Eqs. (15) and (19), we obtain
@@
To have consistency with the AdS/CFT correspondence, $L$ must depend on a positive power of $N$ in the sense of obtaining a low curvature regime when $N$ is large [75]. In this case, we must have $2\alpha\varepsilon-3>0$.
Now, the Eqs. (3), (12) and (23), together give
$\frac{R_{\Diamond}}{L}\sim N^{\frac{2\alpha-3}{6}}.$ (24)
By considering such a result, we note that the condition (9) imposes that $2\alpha-3<0$ in the equation above, and consequently
$\alpha<\frac{3}{2}.$ (25)
In this way, the conditions to have the validity of the holographic principle in the present scenario, with the universe’s causal diamond as the unity of holography, impose a finite upper bound on the exponent $\alpha$, and then on the efficiency of quantum correlations to organize in order to accomplish the emergence of spacetime.
From the Eq (15), it will also imply a saturation limit for the holographic complexity associated with the emergent spacetime, in the situation where one has a fixed quantum correlation budget $N$. It is in agreement with the second law of quantum complexity as stated above.
In summary, in the present scenario, quantum complexity saturation follows from holographic bounds when one considers the universe’s causal diamond as the unity of holography, where the saturation of complexity occurs when the capacity of the causal diamond to accomodate quantum informations reaches its maximum.
By considering the exponent $\gamma$, we have that the only possible solution to it, given by Eq. (22), that is consistent with the existence of an upper bound to the exponent $\alpha$ is
@@
In fact, the results above provide:
$\frac{3}{2\varepsilon}<$ $\alpha$ $\leq\frac{3(\varepsilon+1)}{4\varepsilon}<\frac{3}{2},$ (27)
with
$\varepsilon>1.$ (28)
@@
In this scenario, by considering the second law of holographic complexity, $H_{0}$ must increase during the evolution of the cosmos till it saturates at a maximal value. It must shed some light on the issue of Hubble tension by tracing a way $H_{0}$ can have different values throughout the universe’s evolution.
Before proceeding, we note at first that the factor $log(N^{\gamma})$ in Eq. (21) corresponds to a logarithmic measure of the effective system size, rescaled by the cosmological scaling $\gamma$. To simplify our discussions, we shall fix $N^{\gamma}=e$ in this equation, in a way that
$H_{0}=De^{\alpha}\;.$ (29)
It corresponds to fixing the available complexity phase space of the universe, in a way that the behavior of $H_{0}$ will be modeled entirely by the exponent $\alpha$ in the Eq. (29). In this case, only the efficiency of complexity growth will control the cosmological expansion. It will be important for the analysis presented in the next sections.
6 Evolution of holographic complexity.
In this section, we shall analyze the evolution of the universe’s holographic complexity, by considering a fixed quantum correlation budget, under a given cosmological scale, $N^{\gamma}$. In this way, we shall focus on the evolution of the exponent $\alpha$ in Eq. (15).
In order to link our analysis to the second law of holographic complexity, one can associate the exponent $\alpha$ with a probability weight that indicates how likely it is for high-complexity geometrical states - which are the most likely according to the Second Law - to emerge from quantum correlations.
Now, to apply these ideas to the description of a quantum spacetime given by a superposition of spin network states, each of them characterized by a quantum of area $\Delta$, we note at first that each spin network state is, by its own, a superposition of many quantum geometrical states. Consequently, we shall have not only a simple statistical distribution, but a superposition of many.
In this case, by considering the fluctuation of the parameter $\Delta$ when we change the scale used to define the emergent quantum geometry, in order to relate the exponent $\alpha$ to a probability weight, we must use a superstatistics formulation [80, 81]. In this way, we shall write:
$\alpha(k)=\int_{k}^{\infty}f(k^{\prime})e^{-k^{\prime}\Delta}dk^{\prime}.$ (30)
In the expression above, the steepness parameter $k$ will give the level of coarse-graining of the emergent quantum geometrical states, where large positive values of $k$ strongly suppress larger quantum area eigenvalues, while large negative $k$ values favor them. In the case of $k=0$, there is no suppression. In other words, a larger positive $k$ corresponds to a more sharply-localized, fine-grained behavior. On the other hand, a smaller positive $k$ corresponds to a slower decay: it allows more variance in area contributions, implying a less fine-grained surface structure. Negative $k$ values make coarse-graining effects dominant.
In this way, by taking $f(k)$ as a Gaussion distribution,
$f(k)=Ce^{-ak^{\prime 2}-bk^{\prime}},$ (31)
we obtain
$\alpha(k)$ $=$ $\frac{C}{2}\sqrt{\frac{\pi}{a}}\,e^{\frac{(b+\Delta)^{2}}{4a}}\,\textrm{erfc}\left(\sqrt{a}\left(k+\frac{b+\Delta}{2a}\right)\right)$ (32)
$=$ $C^{\prime}\,\textrm{erfc}\left(\sqrt{a}\left(k+\frac{b+\Delta}{2a}\right)\right),$
where erfc is the complementary error function. In the expression above, $a$, $b$, and $C$ are constants. Moreover, we have introduced
$C^{\prime}=\frac{C}{2}\sqrt{\frac{\pi}{a}}\,e^{\frac{(b+\Delta)^{2}}{4a}},$ (33)
as a constant, for a fixed value of $\Delta$.
From the result in Eq. (32), $\alpha(k)$ will correspond to an S-shaped function (a sigmoid function) of the coarse-graining parameter $k$. The behavior of $\alpha$ is shown in Fig. (2).
Figure 2: The exponent $\alpha$, and $H_{0}(k)$ are plotted against the quantum number $k$, by considering $C^{\prime}=0.5\;ln(\frac{73.04}{67.4}))$, $D=67.4$, $a=1.0$, $b=0.0$, and $\Delta=0.1$. Large positive $k$ values strongly suppress larger area eigenvalues corresponding to a more sharply-localized, fine-grained behavior. On the other hand, large $k$ negative values favor larger area eigenvalues, implying a coarse-grained surface structure.
As we can observe, $\alpha$ approaches its maximal value when $k$ is negative, and coarse-graining effects become dominant. By considering Eq. (27), it must occur when the exponent $\varepsilon$ approaches 1 from above, corresponding to a reduction of the quantum noise in the measurement of the emergent area eigenvalues. In this case, as coarse-graining effects become dominant, they must filter out the quantum noise associated with the emergence of spacetime, enhancing the efficiency of this process.
Consequently, the reduction of microscopic noise related to spacetime emergence, due to coarse-graining effects, guides the increase of the holographic complexity related to the emergent geometry. In such a scenario, the saturation of holographic complexity must occur when coarse-graining eliminates as much noise as possible, thereby reaching the upper limit of the efficiency of spacetime emergence from quantum correlations. This is in line with some results related to holographic coarse-graining presented in the literature [82, 83, 84, 85, 86].
It brings us a very interesting result in the present context: that larger quantum area eigenvalues are shown to be more efficient at processing quantum information toward the emergence of classical geometry. It is in agreement with some ideas on holography in a quantum spacetime where area size is viewed as a measure of the efficiency in processing holographic quantum information [87].
7 Evolution of $H_{0}$ through cosmological history.
Now, let us consider the consequences of the discussions above to the behavior of the Hubble-Lemaître parameter.
As we can observe, from Eq. (29), the results obtained for the exponent $\alpha$ also leads to an S-shaped behavior of $H_{0}$ (Fig. (2)). It can reproduce the measured values by early and late-time observations, by a suitable choice of the constants.
In this case, the early-time value occurs at higher redshifts, when the universe was young, and fine-grained emergent quantum geometrical states are dominant, in a way that higher quantum noise effects are more abundant. On the other hand, the late-time values occur at low redshifts when coarse-grained emergent geometrical states become dominant, reducing microscopic noise to a minimum.
Intermediate redshift values of $H_{0}$ can also be fitted in this scenario. In this way, by taking $k=k(z)$, we have that it must change its sign above and below some critical value of the redshift $z_{c}$. In this case, one can likewise expand
$k(z)$ $=k_{0}(z-z_{c})+\textrm{fluctuation terms}$ (34)
$\approx\;k_{0}(z-z_{c}),$
where it is assumed that $k(z)>0$ for the high-redshift phase, while $k(z)<0$ for the low-redshift phase, for a transition to occur. It corresponds to the Landau theory of phase transitions [88].
Based on such considerations, we have plotted our theoretical result for $H_{0}$ in Fig. (3), by assuming $k_{0}=20$ and $z_{c}=0.4$. The plot shows the fitting of our theoretical result with the key measured values for $H_{0}$. In this way, our model can fit the early, late, and middle-time measurements of $H_{0}$, being in agreement with observations.
[…]
@@
In the present paper, we utilized the quantum gravity scenario introduced in [22, 23] to approach the important issue of the Hubble tension. In this case, it was possible to trace an interesting relationship between $H_{0}$ and the holographic complexity associated with the emergence of cosmological spacetime.
@@
In this case, from the results found in the present paper, $H_{0}$ must increase with the universe’s holographic complexity during the evolution of the cosmos till it saturates. At this point, such a parameter must reach a constant behavior, recovering the description given by the standard cosmological model. This is in line with several results that have pointed out the possibility that $H_{0}$ must vary throughout the cosmological evolution [19, 20, 21, 90, 91].
In front of such results, the difference between the values obtained for $H_{0}$ when one considers early and late-time measurements must be only a testimony that our universe has evolved from a relatively simple state at its early times to a very complex place containing structures like galaxies, stars, quasars, black holes, as well as accelerated expansion. In this sense, it is interesting that black holes have been pointed out as the structures that carry the maximal amount of quantum complexity in nature [92].
Such a quantum gravitational reasoning about the Hubble tension can bring us a change of perspective on cosmological observations by taking quantum complexity as a fundamental concept to describe physical reality, with its second law serving as a guide to the emergence of spacetime from quantum correlations.
In this sense, we must note at first that, in the context of the theory introduced in [22, 23], there is no role for observers (quantum reference frames) in a quantum gravitational context. This is because there is no fundamental distinction between quantum reference frames and quantum correlations in this scenario, where only quantum correlations are led on the table as the fundamental degrees of freedom of the theory.
In this way, our results introduce an observer-independent approach to quantum cosmology, where observer-independence occurs at the fundamental microscopic level, where the cosmos itself is described by quantum correlations structured by the universal law of quantum complexity growth, independent of observers. Consequently, the universe must have a microscopic objective reality based on this law.
On the other hand, observer-dependence can be recovered during the emergence of spacetime, where ”observers” must arise as coarse-grained structures, corresponding to different complexity patterns of quantum correlations. In this way, the idea of observer dependence must be replaced by the concept of observer emergence in the scenario proposed in the present paper.
One can catch sight of this by noting that $C_{h}^{-\gamma}\sim H_{0}^{-1}$ gives a distance scale in cosmology, corresponding to the observable universe’s radius. Consequently, it is possible to understand the different $H_{0}$ values as a manifestation of the universe’s complexity dependence on distance. In this case, the universe’s complexity must decrease when we consider large distance scales.
@@
In this way, different observers in the universe emerge only as different measures of the holographic complexity associated with the emergence of spacetime by considering different regions of the cosmos. In such a scenario, when different observers sample the universe, they inherit the same objective quantum complexity law but can translate it into different effective portrayals of the cosmological parameters depending on their complexity budget. It can open up the range of investigation in quantum cosmology by revealing a new perspective of the world where quantum complexity assumes a central role.
Future explorations in Astrophysics can be done to explore the scenario introduced in the present work by considering, e.g., gravitational waves [93], as future detectors become able to measure the quantum complexity amount carried by this kind of signal.
Acknowledgements
The author acknowledges the anonymous referees for useful comments and suggestions.